OpenStudy (anonymous):

evaluate the limit of sin(3u)/u cos(u) as u approaches 0 plz helpppppp

OpenStudy (jdoe0001):

\(\bf lim_{u \to 0} \cfrac{sin(3u)}{u}cos(u) \quad ?\)

OpenStudy (anonymous):

\[\lim u->0 (\sin(3u))/(u \cos(u))\]

OpenStudy (anonymous):